# Bringing closure to FDR control: beating the e-Benjamini-Hochberg procedure

### v1

@@ Abstract
False discovery rate (FDR) has been a key metric for error control in multiple hypothesis testing, and many methods have developed for FDR control across a diverse cross-section of settings and applications. We develop a closure principle for all FDR controlling procedures, i.e., we provide a characterization based on e-values for all admissible FDR controlling procedures. We leverage this idea to formulate the closed eBH procedure, a (usually strict) improvement over the eBH procedure for FDR control when provided with e-values. We demonstrate the practical performance of closed eBH in simulations.
@@ 1 Introduction
$k^{\textnormal{eBH}}$ $\coloneqq\max\ \left\{k\in[K]:\sum_{i\in[K]}\mathbf{1}\left\{E_{i}\geq\frac{K}{\alpha k}\right\}\geq k\right\}$ (1)
@@
We introduce the ${\overline{\text{eBH}}}$ procedure in Section 2. Then, we formulate our closue principle for all FDR controlling procedures in Section 3. In Section 4, we formulate the closed compound eBH procedure for compound e-values. We describe some extensions and synergies of the ${\overline{\text{eBH}}}$ procedure with existing e-value techniques (i.e., randomization, post-hoc validity, e-merging, boosting). We present simulations showing the improvements of ${\overline{\text{eBH}}}$ over existing eBH methodology in Section 6. Lastly, we discuss prior work in depth in Section 7 and recap our findings in Section 8.
@@ 2 The closed eBH (${\overline{\text{eBH}}}$) procedure
$R^{\overline{\textnormal{eBH}}}$ $\coloneqq\mathcal{R}^{\overline{\textnormal{eBH}}}(\mathbf{E})\coloneqq R_{k^{\overline{\textnormal{eBH}}}}.$ (6)
@@
$\overline{\mathcal{C}}^{\textnormal{SC}}$$\coloneqq\left\{R\subseteq[K]:\min_{i\in R}\ E_{i}\geq K/(\alpha|R|)\right\},$ (7)
@@
for any $R\in\mathcal{C}^{\textnormal{SC}}$ by (11). Thus, we get that $\mathcal{C}^{\textnormal{SC}}\subseteq\overline{\mathcal{C}}$ and also that $R^{\textnormal{eBH}}\in\mathcal{C}$ Since $R^{\overline{\textnormal{eBH}}}\in\underset{R\in\overline{\mathcal{C}}}{\text{argmax}}\ |R|$, we have that $|R^{\textnormal{eBH}}|=k^{\textnormal{eBH}}\leq k^{{\overline{\textnormal{eBH}}}}=|R^{\overline{\textnormal{eBH}}}|$ and consequently $R^{\textnormal{eBH}}=R_{k^{\textnormal{eBH}}}\subseteq R_{k^{\overline{\textnormal{eBH}}}}=R^{\overline{\textnormal{eBH}}}$. Thus, we have shown our desired results. ∎
@@
      return $R^{\overline{\textnormal{eBH}}}=R_{k}$ (discovery set of $k$ largest e-values)
@@
3 A closure principle for FDR control
Theorem 1.
$E_{A}=\frac{\textnormal{FDP}_{A}(R)}{\alpha}.$ (14)
to form a e-collection $(E_{A})_{A\subseteq[K]}$ ensures that in $R\in\overline{\mathcal{C}}$. Further, the only resulting e-closed procedure $\overline{\mathcal{R}}$ is equivalent to $\mathcal{R}$, i.e., $\overline{R}=R$.
@@ 3 A closure principle for FDR control
when we treat $0/0=1$ in the above division. Thus, $R\in\overline{\mathcal{C}}$. As a result, there exists a e-closed procedure which can always choose $\overline{R}\in\overline{\mathcal{C}}$. On the other hand, for every $R^{\prime}\subseteq[K]$ with $|R^{\prime}|>|R|$, we have any $A=R\cap R^{\prime}\cup A^{\prime}$ where $\emptyset\subset A^{\prime}\subseteq R^{\prime}\setminus R$ has $E_{A}=\textnormal{FDP}_{A}(R)/\alpha<\textnormal{FDP}_{A}(R^{\prime})/\alpha$. Similarly, if $|R^{\prime}|<|R|$, we can choose $A^{\prime}\in R\setminus R^{\prime}$ instead and note the same inequality. Thus, $R$ is the only set in $\overline{\mathcal{C}}$.
Remark 1.
The offline multiple testing problem can be viewed as a composite generalized Neyman-Pearson (GNP) problem (Grünwald, 2024). In particular, $R$ is the maximally compatible decision rule for the multiple testing problem w.r.t. the e-collection derived (from $R$) in Theorem 1. However, admissibility results for maximally compatible decision rules from Grünwald (2024) do not directly apply to our setting since their assumption on action “richness” (continuity of loss — in our case, FDP — w.r.t. the action space) is violated in multiple testing, as there are only a finite set of values that the FDP can take.
4 The compound ${\overline{\text{eBH}}}$ procedure

### v2

@@ Abstract
False discovery rate (FDR) has been a key metric for error control in multiple hypothesis testing, and many methods have developed for FDR control across a diverse cross-section of settings and applications. We develop a closure principle for all FDR controlling procedures, i.e., we provide a characterization based on e-values for all admissible FDR controlling procedures. A general version of this closure principle can recover any multiple testing error metric and allows one to choose the error metric post-hoc. We leverage this idea to formulate the closed eBH procedure, a (usually strict) improvement over the eBH procedure for FDR control when provided with e-values. This also yields a closed BY ($\overline{\text{BY}}$) procedure that dominates the Benjamini-Yekutieli (BY) procedure for FDR control with arbitrarily dependent p-values, thus proving that the latter is inadmissibile. We demonstrate the practical performance of our new procedures in simulations.
@@ 1 Introduction
$k^{\textnormal{eBH}}$ $\coloneqq\max\ \left\{k\in\{0\}\cup[K]:\sum_{i\in[K]}\mathbf{1}\left\{E_{i}\geq\frac{K}{\alpha k}\right\}\geq k\right\}$ (1)
@@
Contributions.
Our contributions are listed below.
  • Development of a general closure principle that encompasses all expectation based error rates and allows to construct all procedures controlling these metrics (Theorem 3). This characterizes multiple testing procedures by single e-values for the intersection hypotheses and often yields improvements of existing procedures. While closure principles for FWER control (Marcus et al., 1976) and FDP probability bounds (Goeman and Solari, 2011) are known, our principle additionally includes FDR control which was an open problem, while also recovering these earlier known closure principles.
  • Illustration of the application of our closure principle by introducing a closed eBH (${\overline{\text{eBH}}}$) and a closed BY ($\overline{\text{BY}}$) procedure that dominate the original methods (Section 2). Both eBH and BY are well-studied and were not generally known to be inadmissible (it was known that BY can be improved by randomization, but not by a deterministic procedure (Xu and Ramdas, 2023)).
  • Showing that in its most general form our closure principle allows for post-hoc choice (in a specific sense) of the significance level $\alpha$ (Section 3) and for post-hoc choice of the error metric (Section 4), thus providing multiple testing guarantee that has not been considered before.
@@
We introduce the ${\overline{\text{eBH}}}$ procedure in Section 2. In Section 2.2, we formulate the closed compound eBH procedure for compound e-values. Then, we formulate our closure principle for all FDR controlling procedures in Section 3 and then extend it to any expectation based error rate in Section 4. We describe some extensions and synergies of the ${\overline{\text{eBH}}}$ procedure with existing e-value techniques in Section 5 (i.e., randomization, e-merging, boosting). We present simulations showing the improvements of ${\overline{\text{eBH}}}$ over existing eBH methodology in Section 6. Lastly, we discuss prior work in depth in Section 7 and recap our findings in Section 8.
@@ 2 The closed eBH (${\overline{\text{eBH}}}$) procedure
$R^{\overline{\textnormal{eBH}}}$ $\coloneqq\mathcal{R}^{\overline{\textnormal{eBH}}}(\mathbf{E})\coloneqq R_{[:k^{\overline{\textnormal{eBH}}}]}.$ (6)
@@
$\mathcal{C}^{\textnormal{SC}}$$\coloneqq\left\{R\subseteq[K]:\min_{i\in R}\ E_{i}\geq K/(\alpha|R|)\right\},$ (7)
@@
for any $R\in\mathcal{C}^{\textnormal{SC}}$ by (11). Thus, we get that $\mathcal{C}^{\textnormal{SC}}\subseteq\overline{\mathcal{C}}$ and also that $R^{\textnormal{eBH}}\in\mathcal{C}$ Since $R^{\overline{\textnormal{eBH}}}\in\underset{R\in\overline{\mathcal{C}}}{\text{argmax}}\ |R|$, we have that $|R^{\textnormal{eBH}}|=k^{\textnormal{eBH}}\leq k^{{\overline{\textnormal{eBH}}}}=|R^{\overline{\textnormal{eBH}}}|$ and consequently $R^{\textnormal{eBH}}=R_{[:k^{\textnormal{eBH}}]}\subseteq R_{[:k^{\overline{\textnormal{eBH}}}]}=R^{\overline{\textnormal{eBH}}}$. Thus, we have shown our desired results. ∎
@@
Ignatiadis et al. (2024, Section S3.2) formulates the minimally adaptive eBH procedure where one first tests the global null with the arithmetic mean of the e-values, and only once the global null rejects, the eBH procedure is applied at level $\alpha K/(K-1)$. Formally we can define
$k^{\textnormal{eBH}\text{m}}$ $\coloneqq\max\left\{k\in[K]:E_{[K]}\geq\alpha^{-1}\text{ and }\sum_{i=1}^{K}\mathbf{1}\left\{E_{i}\geq\frac{K-1}{\alpha k}\right\}\geq k\right\}\cup\{0\},$ (13)
$R^{\textnormal{eBH}\text{m}}$ $\coloneqq\mathcal{R}^{\textnormal{eBH}\text{m}}(\mathbf{E})\coloneqq R_{[:k^{\textnormal{eBH}\text{m}}]}.$ (14)
They show that this method dominates eBH as well. This method is, in turn, dominated by the ${\overline{\text{eBH}}}$ procedure.
Proposition 3.
$R^{\textnormal{eBH}\textnormal{m}}\subseteq R^{{\overline{\textnormal{eBH}}}}$.
Proof.
By definition of $R^{\textnormal{eBH}\text{m}}$, we know that if any discoveries are made,
$E_{[K]}\geq\alpha^{-1}=\frac{\textnormal{FDP}_{[K]}(R^{\textnormal{eBH}\text{m}})}{\alpha}.$
Consider the following derivation for arbitrary $A\subset[K]$, i.e, $|A|\leq K-1$, and $R\subseteq[K]$:
$E_{A}$ $=|A|^{-1}\sum\limits_{i\in A}E_{i}$ (15)
$\geq|A|^{-1}|A\cap R|\min_{i\in A\cap R}E_{i}\geq|A|^{-1}|A\cap R|\min_{i\in R}E_{i}\geq(K-1)^{-1}|A\cap R|\min_{i\in R}E_{i}.$ (16)
Recall that $E_{A}$ is defined as the arithmetic mean e-value in (2). The first inequality is lower bounding each term in the sum and also dropping some positive terms. The second inequality follows by strictly expanding the set which we take a minimum over. The third inequality is because $|A|\leq K-1$.
Now, consider $A$ where $\textnormal{FDP}_{A}(R^{\textnormal{eBH}\text{m}})>0$ (otherwise $\textnormal{FDP}_{A}(R^{\textnormal{eBH}\text{m}})/\alpha\leq E_{A}$ is trivially satisfied). We have the following result:
$E_{A}\geq(K-1)^{-1}|A\cap R|\min_{i\in R^{\textnormal{eBH}\text{m}}}E_{i}\geq\frac{|A\cap R^{\textnormal{eBH}\text{m}}|}{\alpha|R^{\textnormal{eBH}\text{m}}|}=\frac{\textnormal{FDP}_{A}(R^{\textnormal{eBH}\text{m}})}{\alpha}.$ (17)
The first inequality is by (16), and the second inequality is by the definition of $R^{\textnormal{eBH}\text{m}}$. Thus, $R^{\textnormal{eBH}\text{m}}\in\overline{\mathcal{C}}$ and we have that $R^{\textnormal{eBH}\text{m}}\subseteq R^{\overline{\textnormal{eBH}}}$. ∎
Where is ${\overline{\textnormal{eBH}}}$ gaining power?
To better illustrate the features of the ${\overline{\textnormal{eBH}}}$ procedure, we can take a close look at the thresholds on the individual e-values for rejection. Choose $A=\{j\}$ for each $j\in R_{[:k]}$, we see that $\overline{\mathcal{C}}$ only enforces the constraint that
$E_{j}\geq\frac{1}{\alpha k^{\overline{\textnormal{eBH}}}}\text{ for every }j\in R^{\overline{\textnormal{eBH}}}.$
Of course, it imposes other “joint constraints” on averages of these e-values as well: for instance, the average of the smallest $m$ rejected e-values must esceed $m/(\alpha k^{{\overline{\textnormal{eBH}}}})$. In contrast, the eBH procedure satisfies
$E_{j}\geq\frac{1}{\alpha}\text{ for every }j\in R^{\textnormal{eBH}},$
which is a significantly harsher threshold. Meanwhile, the minimally adaptive eBH procedure that makes $k^{\textnormal{eBH}\text{m}}\geq k^{\textnormal{eBH}}$ rejections satisfies
$E_{j}\geq\frac{1}{\alpha\tfrac{K}{K-1}}\text{ for every }j\in R^{\textnormal{eBH}\text{m}},$
which is again a much more stringent condition. Since $K/(K-1)\leq 2$, with equality only for $K=2$, the latter threshold matches that of the closed eBH procedure for $K=2$, but is more stringent otherwise. It can be easily shown that the closed eBH procedure indeed equals the minimally adaptive eBH procedure for $K=2$, but can be more powerful otherwise, as demonstrated by the following simple example.
Example 1 (Closed eBH beats minimally adaptive eBH).
We give a simple example with $K=3$ to show that the closed eBH rule can reject strictly more than the minimally adaptive eBH rule. For the set of input e-values $\mathbf{E}=(60,39,11)$, the eBH procedure only makes one rejection at $\alpha=0.05$, narrowly missing out on the second rejection. The minimally adaptive eBH procedure effectively runs eBH at level $23\alpha/$ 2, and makes two rejections. The closed eBH procedure rejects all 3 hypotheses, which can be checked by verifying that the set $\{1,2,3\}$ lies in $\overline{\mathcal{C}}$, since each e-value is larger than 20/3, all pairwise averages exceed 40/3 and the overall average exceeds 20.
@@
      return $R^{\overline{\textnormal{eBH}}}=R_{[:k]}$ (discovery set of $k$ largest e-values)
@@
Post-hoc validity
A stronger form of control on the FDP has been considered in prior literature (Grünwald, 2024; Xu et al., 2024; Koning, 2024) that the eBH (or any e-self-consistent) procedure satisfies. To define this guarantee, we let $\overline{\mathcal{C}}_{\beta}$ denote $\overline{\mathcal{C}}$ as defined in (3) constructed for level $\beta\in[0,1]$.
Proposition 4.
The following post-hoc validity is satisfied for the set of $(\overline{\mathcal{C}}_{\beta})_{\beta\in[0,1]}$.
$\mathbb{E}\left[\sup_{\beta\in[0,1]}\sup_{R\in\overline{\mathcal{C}}_{\beta}}\ \frac{\textnormal{FDP}_{A^{*}}(R)}{\beta}\right]\leq 1,$ (18)
where we let $0/0=1$.
@@
By construction of $\overline{\mathcal{C}}_{\beta}$ for each $\beta\in[0,1]$, we have the following deterministic inequality
$\sup_{\beta\in[0,1]}\sup_{R\in\overline{\mathcal{C}}_{\beta}}\ \frac{\textnormal{FDP}_{A^{*}}(R)}{\beta}\leq E_{A^{*}}.$ (19)
We get our desired result by taking expectations on both sides. ∎
Thus, we get a form of control even when we choose both $\beta$ and $R$ in a data-dependent fashion.
2.1 The closed BY ($\overline{\text{BY}}$) procedure
Our results also have implications for the Benjamini and Yekutieli (2001) (BY) procedure, which ensures FDR control under arbitrary dependence among p-values $\mathbf{X}=\mathbf{P}\coloneqq(P_{1},\dots,P_{K})\in[0,1]^{K}$. The procedure is defined as follows:
$k^{\textnormal{BY}}$ $\coloneqq\max\left\{k\in\{0\}\cup[K]:\sum_{i=1}^{k}\mathbf{1}\left\{P_{i}\leq\frac{\alpha k}{K\ell_{K}}\geq k\right\}\right\}$ (20)
$R^{\textnormal{BY}}$ $\coloneqq\mathcal{R}^{\textnormal{BY}}(\mathbf{P})\coloneqq R_{(:k^{\textnormal{BY}})},$ (21)
where $R_{(:k)}$ denotes the set of hypotheses corresponding to the $k$ smallest p-values.
Define the following p-to-e calibrator $f^{\textnormal{BY}(\alpha,K)}:[0,1]\rightarrow\mathbb{R}^{\geq 0}$ (Xu et al., 2024):
$f^{\textnormal{BY}(\alpha,K)}(x)\coloneqq K\alpha^{-1}\left(\left\lceil x(\alpha/(K\ell_{K}))^{-1}\right\rceil\vee 1\right)^{-1}\mathbf{1}\{x\leq\alpha^{-1}\ell_{K}\},$ (22)
where $\ell_{K}$ is the $K$th harmonic number. Wang and Ramdas (2022) showed that that this is equivalent to applying eBH to
$\mathbf{E}^{\textnormal{BY}(\alpha,K)}=(f^{\textnormal{BY}(\alpha,K)}(P_{1}),\dots,f^{\textnormal{BY}(\alpha,K)}(P_{K})).$
Hence, we can now the $\overline{\text{BY}}$ procedure as
$R^{\overline{\textnormal{BY}}}\coloneqq\mathcal{R}^{\overline{\textnormal{BY}}}(\mathbf{P})\coloneqq\mathcal{R}^{\overline{\textnormal{eBH}}}(\mathbf{E}^{\textnormal{BY}(\alpha,K)}).$ (23)
Proposition 5.
$R^{\textnormal{BY}}\subseteq R^{\overline{\textnormal{BY}}}$.
The above proposition follows directly from the formulation of BY as eBH and Proposition 2. We also see empirically in Section 6, which gives us the rather surprising result that the BY procedure is inadmissible, even when restricted to deterministic procedures (as Xu and Ramdas (2023) provided a randomized version of the BY procedure that dominated the original).
2.2 The compound ${\overline{\text{eBH}}}$ procedure
@@
3 A closure principle for FDR control
Theorem 1.
$E_{A}=\frac{\textnormal{FDP}_{A}(R)}{\alpha}.$ (29)
to form a e-collection $(E_{A})_{A\subseteq[K]}$ ensures that in $R\in\overline{\mathcal{C}}$. Further, the dominant e-closed procedure $\overline{\mathcal{R}}$ is equivalent to $\mathcal{R}$, i.e., $\overline{R}=R$.
Proof.
when we treat $0/0=1$ in the above division. Thus, $R\in\overline{\mathcal{C}}$. As a result, there exists a e-closed procedure which can always choose $\overline{R}\in\overline{\mathcal{C}}$. On the other hand, for every $R^{\prime}\subseteq[K]$ with $|R^{\prime}|>|R|$, we have any $A=R\cap R^{\prime}\cup A^{\prime}$ where $\emptyset\subset A^{\prime}\subseteq R^{\prime}\setminus R$ has $E_{A}=\textnormal{FDP}_{A}(R)/\alpha<\textnormal{FDP}_{A}(R^{\prime})/\alpha$. Similarly, if $0<|R^{\prime}|<|R|$, we can choose $A^{\prime}\in R\setminus R^{\prime}$ instead and note the same inequality. Thus, $R$ and $\emptyset$ are the only sets in $\overline{\mathcal{C}}$. As a result, letting the e-closed procedure always outputting $R$ is the dominant procedure.
Closure for post-hoc valid procedures.
We also formulate an alternative way to form an e-collection for the e-closed procedure from procedures with post-hoc FDP control (in the sense of (18)). Here, let $\mathcal{D}$ be a procedure that produces a family of discovery sets $(R_{\beta})_{\beta\in[0,1]}$, and $\mathcal{D}$ is post-hoc FDP controlling procedure if it satisfies
$\mathbb{E}\left[\sup_{\beta\in[0,1]}\frac{\textnormal{FDP}_{A^{*}}(R_{\beta})}{\beta}\right]\leq 1.$ (31)
We can define the following e-values for each $A\subseteq[K]$:
$\bar{E}_{A}=\sup_{\beta\in[0,1]}\frac{\textnormal{FDP}_{A}(R_{\beta})}{\beta}.$ (32)
Theorem 2.
For an arbitrary post-hoc procedure $\mathcal{D}$, the e-closed procedure $\overline{\mathcal{R}}$ with e-collection $(\bar{E}_{A})_{A\subseteq[K]}$ that ouputs $\overline{R}$ recovers or improves upon the FDR controlling procedure $\mathcal{R}$ where $R=\mathcal{R}(\mathbf{X})=\mathcal{D}(\mathbf{X})_{\alpha}$ for a specific $\alpha\in[0,1]$.
Proof.
The theorem follows from $(\bar{E}_{A})_{A\subseteq[K]}$ being a valid e-collection due to $\mathcal{D}$ having post-hoc FDP control. Further, we note that $\bar{E}_{A}\geq E_{A}$ (as defined in (29)) since $E_{A}$ is included in the supremum taken in $\bar{E}_{A}$. As a result, the resulting e-closed procedure $\overline{R}$ provides FDR control by outputting a discovery set in $\overline{R}\in\overline{\mathcal{C}}$ s.t. $\overline{R}\supseteq R$ — this set always exists since $R\in\overline{\mathcal{C}}$. Thus, we have shown our desired result. ∎
Let us briefly compare the construction of the e-collection in (32) to the one for (29). Notably, in (32), the e-values do not depend upon $\alpha$. Further if eBH is plugged into the e-closed procedure applied to e-values from (29), it cannot be improved. However, every e-value in (32) is at least as large, and typically larger, than the corresponding e-value defined in (29) (as the latter is included in the supremum of the former). As a result, typically, taking the closure actually can increase the power over the original procedure. Further, e-collection of (29) sets $E_{i}=0$ for $i\not\in R$. On the other hand, the e-collection of (32) constructs $E_{i}$ s.t. $E_{i}>0$ is possible for $i\not\in R$, and they are usually positive values. Thus, the e-closed procedure derived from (32) could improve over the FDR controlling procedure $\mathcal{R}$ used to derive the e-collection.
4 Closed testing for general error metrics
Suppose we have a general error metric $\text{F}=(\text{F}_{A})_{A\subseteq[K]}$, which means that $\text{F}_{A}$ is a (possibly random) function that maps any rejection set $R$ to some nonnegative real number and we aim to find an $R$ such that
$\sup_{A\subseteq 2^{[K]}}\mathbb{E}_{A}\left[\text{F}_{A}(R)\right]\leq\alpha.$ (33)
For example, in case of FDR control, we have $\text{F}_{A}=\text{FDP}_{A}$. However, there are many other error rates this general setup includes:
  • k-familywise error rate (k-FWER): $\text{F}_{A}(R)=\mathbf{1}\{|A\cap R|\geq k\}$.
  • Per-family wise error rate (PFER): $\text{F}_{A}(R)=|A\cap R|$.
  • False discovery exceedance (FDX): $\text{F}_{A}(R)=\mathbf{1}\{\textnormal{FDP}_{A}(R)>\gamma\}$ for some $\gamma\in(0,1)$.
Analogous as before, given an e-value $E_{A}$ for each $A\subseteq[K]$, define the set of candidate discovery sets for the error metric F by
$\overline{\mathcal{C}}(F)\coloneqq\left\{R\subseteq[K]:E_{A}\geq\frac{\text{F}_{A}(R)}{\alpha}\quad\forall{A\subseteq[K]}\right\}.$ (34)
The validity and admissibility of this general closure principle follows exactly as in Proposition 1 and Theorem 1, respectively.
Theorem 3.
Any set $R\in\overline{\mathcal{C}}(F)$ satisfies (33). Furthermore, for any rejection set $R$ and $E_{A}=F_{A}(R)/\alpha$, we have $R\in\overline{\mathcal{C}}(F)$. Hence, every procedure can be obtained as a closed procedure.
This general formulation has the further advantage that it allows the post-hoc choice of the error metric. Let $\mathcal{F}$ be the set of all error metrics under consideration and suppose for every $\text{F}\in\mathcal{F}$, we pick a specific closed procedure $\overline{R}_{F}\in\overline{\mathcal{C}}(F)$ (usually the largest set among all candidates). Then, we have
$\sup_{A\subseteq 2^{[K]}}\mathbb{E}_{A}\left[\sup_{\text{F}\in\mathcal{F}}\text{F}_{A}(\overline{R}_{F})\right]\leq\sup_{A\subseteq 2^{[K]}}\mathbb{E}_{A}\left[E_{A}\right]\alpha\leq\alpha.$ (35)
Hence, we can choose the error metric and thus the corresponding rejection set $\overline{R}_{F}$ post-hoc.
Example 2 (Deciding between e-Holm and ${\overline{\textnormal{eBH}}}$ post-hoc).
Let $E_{A}=|A|^{-1}\sum_{i\in A}E_{i}$ be defined as in Section 2. We know that in case of $\text{F}_{A}=\text{FDP}_{A}$, the resulting closed procedure is given by ${\overline{\textnormal{eBH}}}$. Now suppose we seek for FWER control and thus have $\text{F}_{A}(R)=\mathbf{1}\{|A\cap R|>0\}$. Then,
$\overline{\mathcal{C}}(F)$ =$=\left\{R\subseteq[K]:E_{A}\geq\frac{\text{F}_{A}(R)}{\alpha}\quad\forall{A\subseteq[K]}\right\}$ (36)
=$=\left\{R\subseteq[K]:E_{A}\geq\frac{1}{\alpha}\text{ for all }A\text{ with }A\cap R\neq\emptyset\right\}.$ (37)
Therefore, the largest rejection set $\overline{R}\in\overline{\mathcal{C}}(F)$ contains all indices $i\in[K]$ such that $E_{A}\geq 1/\alpha$ for all $A$ with $i\in A$. This is the e-Holm procedure (Vovk and Wang, 2021; Hartog and Lei, 2025). Consequently, we could decide based on the data if we want to control the FDR with the ${\overline{\textnormal{eBH}}}$ procedure or to control the FWER with the e-Holm procedure without violating the guarantee.
Remark 1.
Our general closure principle recovers the existing closure principles for FWER control (Marcus et al., 1976) and simultaneous FDP control (Goeman and Solari, 2011) by choosing $E_{A}=\phi_{A}/\alpha$ for some level-$\alpha$ intersection tests $\phi_{A}$, $A\subseteq[K]$.
To see this, note that in the FWER case, $\text{F}_{A}(R)=\mathbf{1}\{|A\cap R|>0\}$, the largest set $\overline{R}\in\overline{\mathcal{C}}(F)$ contains all indices $i\in[K]$ with $E_{A}\geq\alpha^{-1}\Leftrightarrow\phi_{A}=1$ (see also Example 2), which is precisely the closure principle for FWER control (Marcus et al., 1976).
For simultaneous FDP control, fix a candidate rejection set $R\subseteq[K]$, set $F_{A}^{d}(R)=\mathbf{1}\{|R\setminus A|<d\}$ and define
$\boldsymbol{d}(R)$ $\coloneqq\max\{d\in\{0,\ldots,K\}:\phi_{A}\geq F_{A}^{d}(R)\quad\forall A\subseteq[K]\}$ (38)
$=\min\{|R\setminus A|:A\subseteq[K],\phi_{A}=0\},$ (39)
then $\boldsymbol{d}(R)$ is the closed procedure for simultaneous true discovery guarantee, which is equivalent to simultaneous FDP control (Goeman et al., 2021). By the simultaneous control of our closure principle over all error rates $F^{d}$, $d\in\{0,\ldots,K\}$, and all rejection sets $R\in\overline{\mathcal{C}}(F^{d})$, we obtain the simultaneous true discovery of $\boldsymbol{d}$ using our general closure principle
$\mathbb{E}_{A^{*}}\left[\sup_{d\in\{0,\ldots,K\}}\sup_{R\in\overline{\mathcal{C}}(F^{d})}F_{A^{*}}^{d}(R)\right]\leq\alpha$ (40)
[…]
Remark 2.
The offline multiple testing problem can be viewed as a composite generalized Neyman-Pearson (GNP) problem (Grünwald, 2024) for different loss functions. In particular, we can view a discovery procedure, $\mathcal{R}$ as the maximally compatible decision rule for the multiple testing problem w.r.t. the e-collection derived in Theorem 1 for FDR control. However, admissibility results for maximally compatible decision rules from Grünwald (2024) do not directly apply to our setting since their assumption on action “richness” (continuity of loss — for FDP or other discrete multiple testing error metrics — w.r.t. the action space) is violated in multiple testing, as there are only a finite set of values that the FDP can take.